### v1

# LLMInit: A Free Lunch from Large Language Models for Selective Initialization of Recommendation

Abstract.
Collaborative filtering models, particularly graph-based approaches, have demonstrated strong performance in capturing user-item interactions for recommendation systems. However, they continue to struggle in cold-start and data-sparse scenarios. The emergence of large language models (LLMs) like GPT and LLaMA presents new possibilities for enhancing recommendation performance, especially in cold-start settings. Despite their promise, LLMs pose challenges related to scalability and efficiency due to their high computational demands and limited ability to model complex user-item relationships effectively. In this work, we introduce a novel perspective on leveraging LLMs for CF model initialization. Through experiments, we uncover an embedding collapse issue when scaling CF models to larger embedding dimensions. To effectively harness large-scale LLM embeddings, we propose innovative selective initialization strategies utilizing random, uniform, and variance-based index sampling. Our comprehensive evaluation on multiple real-world datasets demonstrates significant performance gains across various CF models while maintaining a lower computational cost compared to existing LLM-based recommendation approaches.
Collaborative Filtering, Recommendation, Large Language Models
doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
1. Introduction
Recommender systems (RecSys), particularly collaborative filtering (CF), play a pivotal role in many online platforms and web applications  (Koren et al., 2021; Fan et al., 2019; Zou et al., 2024), aiding users in navigating vast information by offering personalized suggestions. Among the various techniques, collaborative filtering (Wang et al., 2019; He et al., 2020; Xu et al., 2023; Zhang et al., 2024a; Xu et al., 2025) stands out for its effectiveness in predicting user preferences by analyzing past user-item interactions. A key principle among these models is learning randomly initialized user/item representations with carefully designed loss functions to capture each user’s preference (Rendle et al., 2012). However, such CF approaches heavily rely on the observed user-item interactions to optimize the CF embeddings from random initialization and perform poorly in real-world applications where interaction data is sparse and users/items are cold-start.
More recently, the rapid advancement of large language models (LLMs) such as GPT-4  (Achiam et al., 2023) and LLaMa  (Touvron et al., 2023), which have exhibited remarkable proficiency in understanding and processing textual information, have piqued the interest of researchers seeking to explore their potential in improving recommender systems beyond traditional approaches  (Geng et al., 2022; Yuan et al., 2023; Li et al., 2023). One of the most promising directions in this exploration involves adapting LLMs as recommender systems through prompt engineering and tailored instruction tuning. Such integration of LLMs enables the RecSys to better understand complex user preferences, and interpret contextual information, thereby surpassing the traditional models in cold-start settings.
Figure 1. An investigation of the embedding collapse issue of CF model LightGCN (He et al., 2020) in two Amazon datasets. Increasing embedding size with exponentially growing parameters finally leads to performance degradation.
Despite the inspiring progress made in LLMs for recommendations, their adoption in online deployments faces significant challenges related to efficiency and scalability. These issues stem from the inherently time-consuming and computationally intensive nature of LLMs. Such problems can be further exaggerated in real-world applications of large-scale users and items due to the enlarged vocabulary tokens and in-context input. Moreover, LLMs fall short of complex CF interaction understanding owing to the next-token prediction pipeline and the constraints of a maximum token limit (e.g., 4096 tokens in LLaMA) to incorporate the whole CF interactions. This limitation hampers their ability to capture and effectively model global user-item dependencies as in CF models, leading to poor performance in the full-ranking (Hou et al., 2024) and warm-start settings (Bao et al., 2023). Therefore, we raise a critical research question:
How can we harness the power of LLMs for recommendation systems in an effective and scalable manner?
To answer the above question, we propose to utilize LLMs for CF-based RecSys from a new perspective by taking pretrained large language models as a free lunch for the embedding initialization of CF models. One trivial way is to directly utilize the LLMs to generate embeddings with substantial textual information. However, storing the large-scale embeddings from LLMs for each user and item is not efficient nor scalable, especially when the platform scales to millions of users and items. Furthermore, as illustrated in Figure 1, unlike the scaling laws observed in NLP models (Kaplan et al., 2020), the performance of the recommendation model LightGCN (He et al., 2020) deteriorates even when the numbers of parameters scale up exponentially. This behavior resembles the embedding collapse issue (Guo et al., [n. d.]) observed in click-through rate prediction models (Guo et al., 2017; Wang et al., 2021). These findings suggest that significantly smaller embedding sizes (e.g., 128) are more suitable for recommendation tasks. In contrast, the top-10 models in the LLM embedding benchmark MTEB  [^1] https://huggingface.co/spaces/mteb/leaderboard (Muennighoff et al., 2022), pretrained for various NLP tasks, use an average large embedding size of 4,506.
In this paper, we propose LLMInit, a novel initialization framework to bridge the embedding gaps between LLMs and CF models for effective and scalable recommendations. Specifically, we propose three strategies for selecting and transferring essential knowledge from LLM embeddings to CF models. For item embedding initialization, we explore the random selection, uniform selection, and variance-based selection to sample embedding index dimensions from LLM-generated representations. On the user side, where contextual information is frequently missing, we aggregate embeddings via a pooling operation over item neighborhoods. These approaches enable small CF models to selectively inherit the rich semantics of large language model embeddings while maintaining efficiency and scalability during the training and inference stage. Extensive experiments demonstrate the efficiency of LLMInit when applied to several state-of-the-art (SOTA) CF models and compared to other LLM-based models. We summarize our contributions as follows:
  • Conceptually, we identify the embedding collapse issue in the CF-based recommendations and propose a new paradigm of effectively leveraging LLMs for CF embedding initialization.
  • Methodologically, we devise three types of selective initialization strategies to inherit the rich world knowledge from pretrained LLMs into lightweight CF models for scalable recommendations.
  • Empirically, we implement LLMInit in a plug-and-play manner on various SOTA CF models, achieving significant performance and efficiency gains, especially compared to LLM-based RecSys.
2. Preliminaries
2.1. Problem Definition
Recommender systems (RecSys) aims to improve the platform’s accuracy in personalization by leveraging content features and user-item interactions. Formally, consider a textual bipartite graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that models user-item interactions with node-level content features. The node set $\mathcal{V}$ is divided into users $u_{i}\in\mathcal{V}_{u}$ and items $v_{j}\in\mathcal{V}_{v}$. Edges $(u_{i},v_{j})\in\mathcal{E}$ represent interactions or relationships between users and items. The objective is to develop a recommendation algorithm that utilizes the node features and graph structure to predict and rank items that a user is likely to be interested in but has not yet interacted with.
Figure 2. An Illustration of LLMInit framework including contextual LLM input, three types of selective item embedding initialization, and the user embedding aggregation.
2.2. Related Work
LLMs for Recommendation. Recent advancements in this domain can be categorized into three paradigms (Wu et al., 2024): extracting embeddings for sequential recommendation (Qiu et al., 2021; Hou et al., 2022), generating semantic tokens to capture user preferences (Li et al., 2023; Xi et al., 2024), and directly adapting LLMs as RecSys through prompting (Hou et al., 2024; Liang et al., 2025) or tuning (Geng et al., 2022). However, they all fail to capture the intricate CF relationship in the user-item bipartite graphs as they either only focus on sequential order patterns (Qiu et al., 2021; Hou et al., 2022) or semantic understandings (Li et al., 2023; Xi et al., 2024; Hou et al., 2024; Zhang et al., 2023; Geng et al., 2022). In addition, they face notable challenges in efficiency and scalability due to the computational demands of LLMs, large-scale embeddings, overloaded vocabulary tokens, and limited in-context inputs in real-world scenarios. In contrast, our work LLMInit, proposes a novel paradigm to leverage LLMs in a scalable manner and addresses ineffective interaction modeling by resorting to CF models.
Table 1. Performance comparison of different CF models on four benchmark datasets in terms of Recall@10 and NDCG@10.
Method Beauty Toys-Games Tools-Home Office-Products
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
LightGCN (He et al., 2020) 0.0910 0.0432 0.0775 0.0360 0.0574 0.0283 0.0745 0.0365
+LLMInit-Rand 0.0960 (+5.5%) 0.0467 (+8.1%) 0.0805 (+3.9%) 0.0387 (+7.50%) 0.0612 (+6.6%) 0.0313 (+10.6%) 0.0773 (+3.8%) 0.0387 (+6.0%)
+LLMInit-Uni 0.1006 (+10.6%) 0.0469 (+8.6%) 0.0806 (+4.0%) 0.0388 (+7.8%) 0.0633 (+10.3%) 0.0319 (+12.7%) 0.0791 (+6.2%) 0.0395 (+8.2%)
+LLMInit-Var 0.1019 (+12.0%) 0.0485 (+12.3%) 0.0808 (+4.3%) 0.0389 (+8.1%) 0.0633 (+10.3%) 0.0317 (+12.0%) 0.0816 (+9.5%) 0.0414 (+13.4%)
SGL (Wu et al., 2021) 0.1017 0.0474 0.0832 0.0380 0.0580 0.0284 0.0669 0.0297
+LLMInit-Rand 0.1069 (+5.1%) 0.0520 (+9.7%) 0.0885 (+6.4%) 0.0418 (+10.0%) 0.0692 (+19.3%) 0.0337 (+18.7%) 0.0810 (+21.1%) 0.0426 (+43.4%)
+LLMInit-Uni 0.1101 (+8.3%) 0.0513 (+8.2%) 0.0920 (+10.6%) 0.0424 (+11.6%) 0.0676 (+16.6%) 0.0333 (+17.3%) 0.0773 (+15.6%) 0.0350 (+17.9%)
+LLMInit-Var 0.1106 (+8.8%) 0.0530 (+11.8%) 0.0927 (+11.4%) 0.0427 (+12.4%) 0.0686 (+18.3%) 0.0339 (+19.4%) 0.0794 (+18.7%) 0.0421 (+41.8%)
SGCL (Zhang et al., 2024b) 0.1027 0.0499 0.0828 0.0382 0.0585 0.0294 0.0647 0.0298
+LLMInit-Rand 0.1094 (+6.5%) 0.0512 (+2.6%) 0.0929 (+12.2%) 0.0418 (+9.4%) 0.0651 (+11.3%) 0.0326 (+10.9%) 0.0770 (+19.0%) 0.0365 (+22.5%)
+LLMInit-Uni 0.1115 (+8.6%) 0.0513 (+2.8%) 0.0923 (+11.5%) 0.0422 (+10.5%) 0.0650 (+11.1%) 0.0327 (+11.2%) 0.0742 (+14.7%) 0.0354 (+18.8%)
+LLMInit-Var 0.1104 (+7.5%) 0.0522 (+4.6%) 0.0941 (+13.6%) 0.0421 (+10.2%) 0.0646 (+10.4%) 0.0327 (+11.2%) 0.0776 (+19.9%) 0.0366 (+22.8%)
3. Method
As in Figure 2 LLMInit, before CF model training, the raw metadata is concatenated and fed to the LLMs. Then the semantic latent embeddings are selectively sampled via one of the CF item embedding initialization approaches followed by user embedding aggregation.
3.1. Selective Item Embedding Initialization
The CF item representation space can be considered a homomorphic projection of the sophisticated world-knowledge representation space in large language models (Sheng et al., 2024). Motivated by this, we propose to distill and selectively utilize the embeddings from LLMs to initialize the embeddings for CF recommendation models.
Formally, $\mathbf{v}\in\mathbb{R}^{N}$ represents the item embedding vector generated by the pre-trained LLMs. Let $K$ denotes the desired embedding dimensionality after selection. Then $\mathcal{I}\subseteq\{0,1,\dots,N-1\}$ is the selected indices. The resulting $K$-dimensional embedding is: $\mathbf{v}_{\text{selected}}=\{\mathbf{v}[i]:i\in\mathcal{I}\}.$ Here, we propose three novel initialization strategies to mitigate the inefficiency and embedding collapse issue for directly adopting the LLMs embeddings as follows.
3.1.1. Uniform Selection (LLMInit-Uni)
The uniform space among selections provides a balanced representation and helps mitigate potential redundancy or over-reliance on the LLM embeddings. By evenly spacing the indices, the strategy ensures that no region of the embedding vector dominates or is neglected across $\mathbf{v}$:
=$\mathcal{I}=\{k\cdot\lfloor N/K\rfloor:k\in\{0,1,\dots,K-1\}\}.$
3.1.2. Random Selection (LLMInit-Rand)
Randomization introduces diversity in the selected dimension patterns. Indices $K$ are randomly sampled from $\{0,1,\dots,N-1\}$ based on the uniform distribution:
$\mathcal{I}\sim\text{Random}(\{0,1,\dots,N-1\}),\quad|\mathcal{I}|=K.$
3.1.3. Variance Selection (LLMInit-Var)
Variance can serve as a heuristic for identifying information-rich and discriminative dimensions in the representation space. By selecting the top-$k$ dimensions of $\mathbf{v}$ with the highest variance across the dataset, this method prioritizes the most distinctive features, improving the separation of potential positive items from candidate pools. Let $\sigma^{2}[j]$ denote the variance of the $j$-th dimension of $\mathbf{h}$ across the dataset. The variance selects the embedding indices as:
$\mathcal{I}=\text{Top-}K(\sigma^{2}[0],\sigma^{2}[1],\dots,\sigma^{2}[N-1]).$
3.2. Aggregated User Embedding Initialization
Based on the item-side embedding, we design an aggregation strategy for user embedding initialization to tackle the privacy situation where the user’s context information is missing. Suppose we observe a user with historical items initial embeddings $\mathbf{v}_{j}$, and we design the aggregation initialization based on the smoothed neighborhood pooling process. The user embedding $\mathbf{u}_{i}$ is finalized as:
(1) $\mathbf{u}_{i}=\sum_{j\in N_{i}}\frac{1}{{|N_{i}|}}\mathbf{v}_{j},$
The normalization employs the degree ${|N_{i}|}$ to temper the magnitude and bias towards popular users after aggregation pooling.
4. Experiments
4.1. Experimental Set Up
Table 2. The statistics of the datasets.
Dataset # Users # Items # Interactions Sparsity
Beauty 10,554 6,087 94,148 99.8534%
Toys-Games 11,269 7,310 95,420 99.8842%
Tools-Home 9,246 6,199 73,250 99.8722%
Office-Products 3,261 1,584 36,544 99.2925%
We follow the 5-core setting in (Zhang et al., 2024b) and experiment on four Amazon  [^2] https://jmcauley.ucsd.edu/data/amazon/links.html real-world datasets: Beauty, Toys-Games, Tools-Home, and Office-Products, as summarized in Table  2. We use a leave-one-out strategy for data splitting and NDCG@10 and Recall@10 metrics for evaluation. For the CF baselines, we select LightGCN (He et al., 2020) as a representative of graph models, SGL (Wu et al., 2021) for self-supervised contrastive learning approaches, and SGCL (Zhang et al., 2024b) for the supervised contrastive learning paradigm. For LLM-based baselines, we use MoRec (Yuan et al., 2023) , LLMRank (Hou et al., 2024), and LLMRec (Wei et al., 2024). For fair comparisons, we set embedding size $K=128$ for all methods and LLM as MPNet (Song et al., 2020) for embedding generation unless specified.
4.2. Performance Comparison
The experimental results demonstrate that the proposed LLMInit strategies significantly enhance performance across all datasets compared to traditional baselines. Among all LLMInit strategies, LLMInit-Var consistently achieves the best results, leveraging variance-based initialization to capture CF and semantic nuances effectively. While LLMInit-Rand and LLMInit-Uni also outperform the baselines, the variance-based strategy emerges as the most robust and adaptable approach. Notably, in the Office-Products dataset, LLMInit delivers an average over 20% boost in NDCG@10 over SGL and SGCL. These findings underscore the robustness and effectiveness of LLMInit strategies, particularly LLMInit-Var, in enhancing recommendation and handling diverse datasets.
4.3. Efficiency Comparison of LLM-based RecSys
Table 3. Comparisons of computation cost and performance with LLM-based RecSys in full-ranking settings.
Methods Beauty Toys-Games
R@10 N@10 # Para. R@10 N@10 # Para.
MoRec (Yuan et al., 2023) 0.0857 0.0397 13M 0.0793 0.0362 14M
LLMRank (Hou et al., 2024) 7B 7B
LLMRec (Wei et al., 2024) 0.0974 0.0472 7B+13M 0.082 0.0389 7B+14M
LLMInit 0.1104 0.0522 0.2M 0.0941 0.0421 0.2M
Table 3 compares various LLM-based RecSys in full-ranking settings, focusing on the computation cost of parameter size and performances on the Beauty and Toys datasets. For the LLMRank (Hou et al., 2024) and LLMRec (Wei et al., 2024), we adopt the LLaMa-7B  [^3] https://huggingface.co/meta-llama/Llama-2-7b model for text generation. MoRec(Yuan et al., 2023), a modality-based model, leverages LLMs to generate textual embeddings. However, its performance is hindered by the use of a unified transform layer for embedding mapping to each item. Additionally, the parameter size for LLM-generated embeddings, initially around 13M/14M, can scale to billions as the total number of users and items surpasses 10 million, posing significant computational and storage challenges. Note that, LLMRank (Hou et al., 2024) directly employs an LLM for prompting but struggles with an incomplete long-context candidate pool and hallucination problem, leading to unreported performance results. LLMRec incorporates LLM for graph augmentation, improving recall to 0.0974 (Beauty) and 0.082 (Toys) but at a much higher computational cost (7B+ parameters). LLMInit, a lightweight alternative, achieves the best performance with only 0.2M parameters, highlighting the potential of parameter-efficient strategies in LLM-powered RecSys.
4.4. Cold-Start Recommendation
Figure 3. Performance improvement comparison in non-strict cold-start scenarios on Amazon-Beauty, where half of the observed interactions for each user are removed.
In the long-tail cold-start setting (Zhang et al., 2025), we remove half of the observed interactions of each user in training, resulting in many cold-start users with only a single interaction. Figure 3 highlights how the three CF models—LightGCN, SGL, and SGCL— react to the limited interactions and benefit from LLM-based initialization strategies in cold-start scenarios. LightGCN, as a simple graph model, is most resistant to the cold-start challenge, while the SGL and SGCL both suffer sharp performance drops due to reliance on interactions for data augmentation or supervised loss optimization. With the LLMInit, SGL demonstrates a stronger ability to utilize enhanced embeddings, by extending LightGCN with self-supervised loss. SGCL, which introduces supervised contrastive learning loss, achieves the highest gains with LLMInit strategies, with Recall and NDCG improving by over 50.4% and 58.9%, respectively. The supervised loss in SGCL amplifies the impact of informative embeddings, allowing it to extract more meaningful patterns from the selected dimensions. Overall, the results reveal that more advanced models like SGCL with nuanced supervised graph contrastive loss are better equipped to exploit high-quality LLM-based initializations.
4.5. The impacts of Large Language Models
Table 4. The Performance of LightGCN on Amazon-Beauty with initialization of different LLMs.
LLMs MPNet4441 gte-Qwen5552 GPT-S6663 GPT-L7773 Stella8884
Embedding Size 768 1,536 1,536 3,072 8,096
Rand R@10 0.0960 0.0936 0.0967 0.0925 0.0996
N@10 0.0467 0.0418 0.0463 0.044 0.0447
Uni R@10 0.1006 0.096 0.098 0.0942 0.0949
N@10 0.0469 0.0428 0.0465 0.0445 0.0417
Var R@10 0.1019 0.0932 0.098 0.096 0.0935
N@10 0.0485 0.0429 0.0465 0.046 0.0412
Table 4 illustrates the performance of LightGCN with different LLMs and initialization strategies (Rand, Uni, Var) on Amazon-Beauty. It observed that all LLMs can enhance LightGCN significantly with proper initialization strategies. Variance-based initialization (Var) generally outperforms random (Rand) and uniform (Uni) strategies by prioritizing high-variance dimensions, capturing the most informative features. Notably, MPNet, with the smallest embedding size (768), achieves the highest Recall@10 (0.1019) and NDCG@10 (0.0485), highlighting the importance of embedding quality over size. While larger models like GPT-L and Stella improve LightGCN with LLMInit, they still lag behind MPNet and GPT-S, emphasizing that LLM alignment with the dataset and strategic initialization matter more than embedding size alone.
5. Conclusion
In this work, we propose LLMInit, an effective and scalable framework that leverages pre-trained large language model embeddings for collaborative filtering model initialization. By employing selective strategies for item embeddings and aggregation pooling for user embeddings, LLMInit addresses data sparsity and cold-start challenges while ensuring scalability and efficiency with low-resource embedding size. Extensive experiments show LLMIint consistently enhancing diverse SOTA CF models, demonstrating LLMs’ new potential research direction to advance recommender systems.